OpenStudy (anonymous):

The specific heat of a certain type of metal is 0.128 J/(g·°C). What is the final temperature if 305 J of heat is added to 80.3 g of this metal initially at 20.0 °C?

OpenStudy (anonymous):

Q=m*cp*delta T --> delta T = Q/ (m*cp)

OpenStudy (anonymous):

and delta T = 29,67 so final temperature is 20+29,67 = 49,67 °C